G issued $4,000,000 of 6%, 3-year convertible bonds on 07-01-14 when the market rate for similar...

  1. G issued $4,000,000 of 6%, 3-year convertible bonds on 07-01-14 when the market rate for similar bonds was 6.5%. The bonds were dated 07-01-14 with interest payable January 01 and July 01. G incurred and paid $47,000 of bond issuance costs. On 07-01-16 after making all of its interest payments, 50% of the bonds were converted into 8,000 shares of G’s $1 par value common stock. G only prepares AJEs every December 31. Prepare the entries G should make on:
  1. 07-01-14
  2. 12-31-14
  3. 01-01-15
  4. 07-01-15
  5. 12-31-15
  6. 01-01-16
  7. 07-01-16
  8. 12-31-16
  9. 01-01-17
  10. 07-01-17
Answer #1
Bond issue price= cash int. * PVIFA(3.25%,6) + Par value * PVIF(3.25%,6)
Bond issue price= (4000000*3%)*5.3726 + 4000000*0.8254
Bond issue price= 3946312
Discount amount = 4000000-3946312 = 53688
Amortisation schedule:
Period Cash interest Interest exp. Dis. amortised Carrying Value
1-Jul-14 3946312
1-Jan-15 120000 128255 8255 3954567
1-Jul-15 120000 128523 8523 3963091
1-Jan-16 120000 128790 8790 3971881
1-Jul-16 120000 129076 9076 3980957
1-Jan-17 120000 129371 9371 3990328
1-Jul-17 120000 129672 9672 4000000
53688
Journal Entries:
Date Acc Titles Debit $ Credit $
7/1/2014 Cash 3946312
Discount on Bond payable 53688
Bond Payable 4000000
(bond issued on dicount)
12/31/2014 Interest expense 128255
Interest payable 120000
Discount on Bond payable 8255
(interest made payable for six months)
1/1/2015 Interest payable 120000
Cash 120000
(interest paid)
7/1/2015 Interest expense 128523
Cash 120000
Discount on Bond payable 8523
(interest paid for six months)
12/31/2015 Interest expense 128790
Interest payable 120000
Discount on Bond payable 8790
(interest made payable for six months)
1/1/2016 Interest payable 120000
Cash 120000
(interest paid)
7/1/2016 Interest expense 129076
Cash 120000
Discount on Bond payable 9076
(interest paid for six months)
7/1/2016 Bond Payable 2000000
Discount on Bond payable 9521 (53688-34645)/2
Common Stock 8000 8000*1
Additional paid in Capital 1982479
(50% bonds converted into common shares and
additional paid-in booked)
12/31/2016 Interest expense 64686
Interest payable 60000
Discount on Bond payable 4686
(interest made payable for six months)
1/1/2017 Interest payable 60000
Cash 60000
(interest paid)
7/1/2017 Bond Payable 2000000
Interest expense 64836
Cash 2060000
Discount on Bond payable 4836
(Principle and interest paid for balance Bonds)